OpenStudy (anonymous):

Someone help me.......... C/5.3+8.3=11.3

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

C/5.3+8.3=11.3 C/5.3=11.3-8.3 C=5.3*(11.3-8.3)

OpenStudy (anonymous):

What?